# Environment-Driven Online LiDAR-Camera Extrinsic Calibration

### v1

@@ III-C Spatial-Temporal Relative Pose Optimization
Multi-Scene Optimization. In each scenario, we choose a reliable subset , which can be obtained as follows:
@@
across multiple spaces in the environment at different times. This process effectively combines optimal correspondences from both spatial and textural pathways in each scenario, enabling environment-driven spatial-temporal optimization and achieving human-like adaptability to the environment.
@@ IV-A Experimental Setup and Evaluation Metrics
In our experiments, we compare our proposed EdO-LCEC with SoTA online, target-free LCEC approaches on the public dataset KITTI odometry [26] (including 00-09 sequences) and MIAS-LCEC [5] (including target-free datasets MIAS-LCEC-TF70 and MIAS-LCEC-TF360). To comprehensively evaluate calibration accuracy, we quantify LCEC performance using the Euler angle error magnitude and the translation error magnitude .
Notably, sequences in KITTI odometry, aside from 00, were included in the training datasets for the learning-based methods [21, 25, 20, 23]. To ensure a fair comparison, we reproduced calibration results for both the left and right cameras on sequence 00 when the authors provided their code; otherwise, we used the reported results for the left camera from their papers. Since most learning-based methods lack APIs for custom data, our comparison with these methods is limited to the KITTI odometry 00 sequence. For experiments on the MIAS-LCEC dataset, as the results for the compared methods [16, 14, 27, 5, 15] are reported in [5], we directly use the values presented in that paper.
@@ IV-B Comparison with State-of-the-Art Method
In this section, quantitative comparisons with SoTA approaches on three datasets are presented in Fig. 5, Tables I222The results generated by LCCNet using the authors’ code yield higher errors compared to those reported in their paper., II, III, and IV. Additionally, qualitative results are illustrated in Figs. 4 and 3.
@@
Evaluation on KITTI odometry. The results shown in Table I and II suggest that, with the exception of sequences 01 and 04, our method achieves SoTA performance across the ten sequences (00-09) in KITTI odometry. Specifically, in the 00 sequence, EdO-LCEC reduces the by around 35.2-99.8% and the by 11.8-98.7% for the left camera, and reduces the by around 46.9-99.7% and the by 13.9-97.6% for the right camera. Additionally, according to Fig. 3, it can be observed that the point cloud of a single frame in KITTI is so sparse that the other approaches behave poorly. In contrast, our proposed method overcomes this difficulty and achieves high-quality data fusion through the calibration result. We attribute these performance improvements to our spatial-temporal relative pose optimization. Merging the optimal matching results from multiple views and scenes maximizes the number of reliable correspondences and ultimately improves overall calibration accuracy.
Evaluation on MIAS-LCEC Dataset. Compared to the sparse point clouds in the KITTI dataset, the point clouds in the MIAS-LCEC datasets are significantly denser, which facilitates feature matching and allows us to test the upper limits of the algorithm calibration accuracy. The results shown in Table III demonstrate that our method outperforms all other SoTA approaches on MIAS-LCEC-TF70. It can also be observed that our method dramatically outperforms CRLF, UMich, DVL, HKU-Mars, and is slightly better than MIAS-LCEC across the total six subsets. In challenging conditions that are under poor illumination and adverse weather, or when few geometric features are detectable, EdO-LCEC performs significantly better than all methods, particularly. This impressive performance can be attributed to the generalizable scene discriminator. The multiple virtual cameras generated by the scene discriminator provide a comprehensive perception of the calibration scene from both spatial and textural perspectives, which largely increases the possibility of capturing high-quality correspondences for the PnP solver. Furthermore, the data fusion results in Fig. 4, obtained using our optimized extrinsic matrix, visually demonstrate perfect alignment on the checkerboard. This highlights the high calibration accuracy achieved by our method.
@@
To evaluate the algorithm’s adaptability to incomplete and sparse point clouds, we further segmented the already limited field-of-view point clouds from the MIAS-LCEC-TF360 dataset. Specifically, as shown in Fig. 2, we divided each point cloud into six equal parts. By progressively combining these segments, we create point clouds with varying densities, ranging from 1/6 (c1) to the full 6/6 (c6) density. The calibration results presented in Fig. 5 show that EdO-LCEC achieves the smallest mean and , along with the narrowest interquartile range, compared to other approaches across different point cloud densities. This demonstrates the stability and adaptability of EdO-LCEC under challenging conditions involving sparse or incomplete point clouds.
IV-C Ablation Study and Analysis
To explore the contribution of scene discriminator and spatial-temporal relative pose optimization, we evaluate our algorithm’s performance with and without multi-view (including both textural and spatial perspective views) and multi-scene optimizations on the 00 sequence of KITTI odometry. The results presented in Table V demonstrate that lacking any of these components significantly degrades calibration performance. In particular, calibration errors increase substantially when only single-view calibration is applied, as it lacks the comprehensive constraints provided by multi-view inputs. Additionally, the joint optimization from multiple scenes significantly improved the calibration accuracy compared to those only under multi-view optimization. These results confirm the advantage of incorporating spatial and textural constraints from multiple views and scenes, validating the robustness and adaptability of our environment-driven calibration strategy.
@@ V Conclusion
In this paper, we explore to extend a new definition called “environment-driven” for online LiDAR-camera extrinsic calibration. Unlike previous methods, our approach demonstrates human-like adaptability across diverse and complex environments. Inspired by human perceptual systems, we designed a scene discriminator that actively analyzes the calibration scenes from spatial and textural perspectives. By leveraging structural and textural consistency between LiDAR projections and camera images, our method achieves reliable 3D-2D correspondence matching, overcoming the challenge of cross-modal feature matching in sparse scenes encountered by earlier approaches. Additionally, we modeled the calibration process as a spatial-temporal joint optimization problem, achieving high-precision and robust extrinsic matrix estimation through multi-view optimization within individual scenes and joint optimization across multiple scenarios. Extensive experiments on real-world datasets demonstrate that our environment-driven calibration strategy achieves state-of-the-art performance.
@@ VI Details of the Algorithm Implementation
Require:
Spatial pathway: Masks (including detected corner points) obtained from the LIP and RGB image.
Textural pathway: Masks (including detected corner points) obtained from the LDP and depth images.
Stage 1 (Reliable mask matching):
(2) Estimate and using (13)-(16).
Stage 2 (Dense correspondence matching):
(1) For each pathway, update all masks in the virtual image using and .
(2) Construct the corner point cost matrix using (6).
(4) Aggregate all corner point correspondences to form the sets .
@@
The pseudo code of DPCM is shown in Algorithm 1. Thanks to the generalizable discriminator, the multiple views generated by the virtual cameras provide comprehensive spatial and textural cues to formulate the feature matching to a 3D-2D correspondence matching process. The entire DPCM is efficient because it primarily focuses on 2D structural and textural consistency computation, rather than complex 3D point cloud registration.
@@ VIII Details of Experimental Configuration
| Borer et al.[20] | WACV 2024 | Learning-based | CNN, Geometric Mutual Information | High | High | No |
@@ VIII-A Datasets
MIAS-LCEC-TF360: MIAS-LCEC-TF360 contains 12 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Since the Livox Mid-360 LiDAR has a scanning range of 360∘, it produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. Additionally, the significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios
VIII-B Metrics
As stated in Sect. IV-A, to comprehensively evaluate the performance of LCEC approaches, we use the magnitude of Euler angle error and the magnitude of the translation error to quantify the calibration errors. Here, we give the specific calculation formula of and 333The translation from LiDAR pose to camera pose is when (1) is used to depict the point translation.:
where and represent the estimated and ground-truth Euler angle vectors, computed from the rotation matrices and , respectively. Similarly, and denote the estimated and ground-truth translation vectors from LiDAR to camera, respectively.
VIII-C Data Preparation
The dataset statistics of KITTI odometry are presented in Table VI. It can be observed that there are a total of 24500 pairs of point clouds and images in KITTI odometry. Note that most of the SoTA comparison methods are slow to complete calibration with a single frame. For the comparison algorithms that provided source code, we performed calibration every ten frames (except for 2 frames in 04 sequence), enhancing the efficiency of the experiments without compromising the validity of the evaluation.
VIII-D Details of Comparison Methods
@@
IX Additional Experiments
IX-A Comparison of Correspondence Matching
IX-B More Visualization Results
In this section, we provide more detailed visualization results of EdO-LCEC through LiDAR and camera data fusion on MIAS-LCEC datasets. Fig. 8 demonstrate that LiDAR point cloud and camera image aligned almost perfectly along the textural and geometric edges, proving the accuracy of our calibration results. It can also be observed that our method performs well in challenging conditions, particularly under poor illumination and adverse weather, or when few geometric features are detectable. This impressive data fusion result further validates the critical role of LiDAR-camera extrinsic calibration. The seamless integration of spatial and textural modalities equips robots with a human-like perceptual ability to interpret and interact with their surroundings. Our environment-driven approach bridges the gap toward the flexible and adaptive machines often envisioned in science fiction.

### v2

@@ III-C Multi-View and Multi-Scene Joint Optimization
In multi-scene optimization, we choose a reliable subset from each scenario, which can be obtained as follows:
@@
across multiple spaces in the environment at different times. This process combines optimal correspondences from both spatial and textural pathways in each scenario, enabling robust environment-driven optimization through multi-view and multi-scene fusion. As a result, our method achieves human-like adaptability to dynamic environments. This environment perception make EdO-LCEC behave much better than our previous method MIAS-LCEC, especially in conditions when point clouds are sparse or incomplete.
@@ IV-A Experimental Setup and Evaluation Metrics
In our experiments, we compare our proposed EdO-LCEC with SoTA online, target-free LCEC approaches on the public dataset KITTI odometry [27] (including 00-09 sequences) and MIAS-LCEC [6] (including target-free datasets MIAS-LCEC-TF70 and MIAS-LCEC-TF360). To comprehensively evaluate the performance of LCEC approaches, we use the magnitude of Euler angle error and the magnitude 222The translation from LiDAR pose to camera pose is when (1) is used to depict the point translation. of the translation error, with the following expression:
to quantify the calibration errors. In (11), and represent the estimated and ground-truth Euler angle vectors, computed from the rotation matrices and , respectively. Similarly, and denote the estimated and ground-truth translation vectors from LiDAR to camera, respectively.
Notably, sequences in KITTI odometry, aside from 00, were included in the training datasets for the learning-based methods [22, 26, 24]. To ensure a fair comparison, we reproduced calibration results for both the left and right cameras on sequence 00 when the authors provided their code; otherwise, we used the reported results for the left camera from their papers. Since most learning-based methods lack APIs for custom data, our comparison with these methods is limited to the KITTI odometry 00 sequence. For experiments on the MIAS-LCEC dataset, as the results for the compared methods [38, 16, 29, 6, 17] are reported in [6], we directly use the values presented in that paper.
Our algorithm was implemented on an Intel i7-14700K CPU and an NVIDIA RTX4070Ti Super GPU. The entire process of a single-view calibration, including scene discriminating, DPCM, and relative pose optimization, takes approximately 15 to 70 seconds.
@@ IV-B Comparison with State-of-the-Art Method
In this section, quantitative comparisons with SoTA approaches on three datasets are presented in Fig. 7, Tables I333The results generated by LCCNet using the authors’ code yield higher errors compared to those reported in their paper., II, III, and IV. Additionally, qualitative results are illustrated in Figs. 4 and 5.
IV-B1 Evaluation on KITTI Odometry
The results shown in Table I and II suggest that, with the exception of sequences 01 and 04, our method achieves SoTA performance across the ten sequences (00-09) in KITTI odometry. Specifically, in the 00 sequence, EdO-LCEC reduces the by around 35.2-99.8% and the by 11.8-98.7% for the left camera, and reduces the by around 46.9-99.7% and the by 13.9-97.6% for the right camera. Additionally, according to Fig. 5, it can be observed that the point cloud of a single frame in KITTI is so sparse that the other approaches behave poorly. In contrast, our proposed method overcomes this difficulty and achieves high-quality data fusion through the calibration result. We attribute these performance improvements to our multi-view and multi-scene joint optimization. Merging the optimal matching results from multiple views and scenes maximizes the number of reliable correspondences and ultimately improves overall calibration accuracy.
IV-B2 Evaluation on MIAS-LCEC Dataset
Compared to the sparse point clouds in KITTI odometry, the point clouds in the MIAS-LCEC datasets are significantly denser, which facilitates feature matching and allows us to test the upper limits of the algorithm calibration accuracy. The results shown in Table III demonstrate that our method outperforms all other SoTA approaches on MIAS-LCEC-TF70. It can also be observed that our method dramatically outperforms CRLF, UMich, DVL, HKU-Mars, and is slightly better than MIAS-LCEC across the total six subsets. In challenging conditions that are under poor illumination and adverse weather, or when few geometric features are detectable, EdO-LCEC performs significantly better than all methods, particularly. This impressive performance can be attributed to the generalizable scene discriminator. The multiple virtual cameras generated by the scene discriminator provide a comprehensive perception of the calibration scene from both spatial and textural perspectives, which largely increases the possibility of capturing high-quality correspondences for the PnP solver. Furthermore, the data fusion results in Fig. 4, obtained using our optimized extrinsic matrix, visually demonstrate perfect alignment on the checkerboard. This highlights the high calibration accuracy achieved by our method.
@@ IV-B2 Evaluation on MIAS-LCEC Dataset
To evaluate the algorithm’s adaptability to incomplete and sparse point clouds, we further segmented the already limited field-of-view point clouds from the MIAS-LCEC-TF360 dataset. Specifically, as shown in Fig. 6, we divide each point cloud into six equal parts based on the recorded temporal sequence of the points. By progressively combining these segments, we create point clouds with varying densities, ranging from 1/6 (c1) to the full 6/6 (c6) density. The calibration results presented in Fig. 7 show that EdO-LCEC achieves the smallest mean and , along with the narrowest interquartile range, compared to other approaches across different point cloud densities. This demonstrates the stability and adaptability of EdO-LCEC under challenging conditions involving sparse or incomplete point clouds.
IV-C Ablation Study and Analysis
| Consistency | Left Camera | Right Camera | |||
| Structural | Textural | (∘) | (m) | (∘) | (m) |
| 0.360 | 0.114 | 0.476 | 0.185 | ||
| ✓ | 0.340 | 0.102 | 0.473 | 0.152 | |
| ✓ | 0.328 | 0.109 | 0.435 | 0.146 | |
| ✓ | ✓ | 0.295 | 0.082 | 0.336 | 0.118 |
@@ IV-C Ablation Study and Analysis
To validate the contribution of structural consistency and textural consistency in DPCM, we conducted an ablation study comparing their individual and combined effects. As shown in Table V, both components significantly improve calibration accuracy. Specifically, structural consistency preserves local geometric features through mask-guided alignment, while textural consistency evaluates visual similarity around matched correspondences. Their combination enables DPCM to maintain robust matching performance even in challenging scenarios with sparse or occluded point clouds, which is a capability that existing methods (DVL and MIAS-LCEC) struggle to achieve. Additionally, MIAS-LCEC only compares the corner points within the matched segmented mask, while EdO-LCEC not only considers the corner points of the matched mask but also includes those of other unmatched masks. This allows EdO-LCEC to achieve global attention when searching the potential available matches when LiDAR frames are sparse.
To explore the contribution of the generalizable scene discriminator as well as the multi-view and multi-scene joint optimization, we evaluate our algorithm’s performance with and without multi-view (including both textural and spatial perspective views) and multi-scene optimizations on the 00 sequence of KITTI odometry. The results presented in Table VI demonstrate that lacking any of these components significantly degrades calibration performance. In particular, calibration errors increase substantially when only single-view calibration is applied, as it lacks the comprehensive constraints provided by multi-view inputs. Additionally, the joint optimization from multiple scenes significantly improved the calibration accuracy compared to those only under multi-view optimization. These results confirm the advantage of incorporating spatial and textural constraints from multiple views and scenes, validating the robustness and adaptability of our environment-driven calibration strategy.
IV-D Comparison of Correspondence Matching
@@ V Conclusion
In this article, we explore extending a new definition called “environment-driven” for online LiDAR-camera extrinsic calibration. Unlike previous methods, our approach introduces environmental observation to maintain optimal performance across diverse and complex conditions. Specifically, we designed a scene discriminator that can automatically observe the calibration scene. This discriminator can guide cross-modal feature matching by evaluating the feature density through multi-modal features extracted by large vision models. By leveraging structural and textural consistency between LiDAR projections and camera images, our method achieves more reliable 3D-2D correspondence matching. Additionally, we modeled the calibration process as a multi-view and multi-scene joint optimization problem, achieving high-precision and robust extrinsic matrix estimation through multi-view optimization within individual scenes and joint optimization across multiple scenarios. Extensive experiments on real-world datasets demonstrate that our environment-driven calibration strategy achieves the state-of-the-art performance.
@@ VIII Details of Experimental Configuration
| Borer et al. [21] | WACV 2024 | MI-based | Geometric Mutual Information | High | High | No |
@@ VIII-A Datasets
MIAS-LCEC-TF360: MIAS-LCEC-TF360 contains 12 pairs of 4D point clouds and RGB images, collected using a Livox Mid-360 LiDAR and a MindVision SUA202GC camera from both indoor and outdoor environments. Since the Livox Mid-360 LiDAR has a scanning range of 360∘, it produces a sparser point cloud compared to that generated by the Livox Mid-70 LiDAR. Additionally, the significant difference in the FoV between this type of LiDAR and the camera results in only a small overlap in the collected data. Consequently, this dataset is particularly well-suited for evaluating the adaptability of algorithms to challenging scenarios.
VIII-B Data Preparation
The dataset statistics of KITTI odometry are presented in Table VIII. It can be observed that there are a total of 24500 pairs of point clouds and images in KITTI odometry. Note that most of the SoTA comparison methods are slow to complete calibration with a single frame. For the comparison algorithms that provided source code, we performed calibration every ten frames (except for 2 frames in 04 sequence), enhancing the efficiency of the experiments without compromising the validity of the evaluation.
The data preparation pipeline for our MIAS-LCEC datasets is illustrated in Fig. 11. Our experimental platform integrates two DJI Livox solid-state LiDARs (Mid-70 and Mid-360) and a MindVision MV-SUA202GC global-shutter CMOS camera for synchronized cross-modal data acquisition. Both LiDARs operate at 200,000 points/s. However, the Mid-70 captures dual returns while the Mid-360 records only the first return. For each scene, we accumulate 15-20 seconds of LiDAR scans to ensure dense point cloud coverage, paired with RGB images at 1,200800 pixel resolution captured by the camera, whose intrinsic parameters were pre-calibrated offline and remained fixed throughout all experiments.
VIII-C Details of Comparison Methods
@@
IX More Visualization Results
In this section, we provide more detailed visualization results of EdO-LCEC through LiDAR and camera data fusion on MIAS-LCEC datasets. Fig. 10 demonstrates that the LiDAR point cloud and camera image aligned almost perfectly along the textural and geometric edges, proving the accuracy of our calibration results. It can also be observed that our method performs well in challenging conditions, particularly under poor illumination and adverse weather, or when few geometric features are detectable. This impressive data fusion result further validates the critical role of LiDAR-camera extrinsic calibration. The seamless integration of spatial and textural modalities equips robots with a human-like perceptual ability to interpret and interact with their surroundings. Our environment-driven approach bridges the gap toward the flexible and adaptive machines often envisioned in science fiction.